Dale invested his savings in two investment funds. The $6000 that he invested in Fund A returned a 10% profit. The amount that he invested in Fund B returned a 3% profit. How much did he invest in Fund B, if both funds together returned a 5% profit?
Answer by richwmiller(17219)

We know 3 rates and one amount invested.
0.1*6000+0.03*x=0.05*(6000+x)
Multiply out.
600+0.03*x=300+0.05*x
Combine like terms.
-0.02*x=-300
Isolate x.
x=-300/-0.02
Divide.
x=$15000
Check.
0.1*6000+0.03(x)=0.05*(6000+x)
600+0.03*15000=0.05*(6000+15000)
600+450=0.05*(21000)
1050=1050
Ok